OpenStudy (anonymous):

Solve the proportion 2x/7x-6 = 2/x.. What is the value of x? x=-1 and x=-7 x=1 and x=-7 x=1 and x=6 x=-1 and x=6

OpenStudy (anonymous):

\[\frac{ 2x }{ 7x-6 } = \frac{ 2 }{ x }\]

OpenStudy (anonymous):